OpenStudy (lexber):

(MEDAL+FAN) 15√49 in simplest form please

OpenStudy (er.mohd.amir):

\[\sqrt{x*x}=x\] AND 49=7*7 USE THIS

OpenStudy (lexber):

@Er.Mohd.AMIR I'm sorry but i don't understand how

OpenStudy (lexber):

the answer choices are fractions A. 7^1/5 B. 7^8/15 C. 7^5/3 D. 7^2/15

OpenStudy (er.mohd.amir):

\[\sqrt{49}=\sqrt{7*7}=7 SO 15\sqrt{49}=?\]

OpenStudy (lexber):

\[14\sqrt{49}\] ?

OpenStudy (anonymous):

15*7=105